Grade 12th passGeneral Physics

A spring of constant K is cut into two parts by the ratio of length 1:2 . The spring constant of larger one is?

Muskan Fatima
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
Dear Student,
From the given data, proportion of spring constants of the two spring sections = 2:1 
What's more, we likewise know the way that the bigger section will have littler spring steady. 
So for this , we can expect the qualities to be 2x and x. At the point when these two are in arrangement, they offer ascent to the first spring. So 
2x * x/(2x + x) = k 
which gives x = 3k/2. 
So spring constants are 3k/2 and 3k. 
When they are in parallel, successful spring consistent = 3k + 3k/2 = 9k/2 
Era T = 2*pi sqrt (2m/9k).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)